x=-4y+4
5x-3y=-3 solve the system by the substitution method

Answer by sprolden(40)

x=-4y+4 solve the system by the substitution method
5x-3y=-3
1- Substitute -4y+4 for x in the bottom equation.
5(-4y+4)-3y=-3
2- Multiply 5 times all terms in parenthesis.
-20y+20-3y=-3
3- Combine like terms.
-20y-3y+20=-3
-23y+20=-3
4- Subtract 20 from each side.
-23y=-23
y=1
5- Substitute 1 into either of the original equations. I'm going to use the first equation.
x=-4(1)+4
x=-4+4
x=0
6- The solutions are x=0 and y=1.
CHECK: 0=-4(1)+4 so 0=0